### v1

# WorldPrompter: Traversable Text-to-Scene Generation

Abstract
Scene-level 3D generation is a challenging research topic, with most existing methods generating only partial scenes and offering limited navigational freedom. We introduce WorldPrompter, a novel generative pipeline for synthesizing traversable 3D scenes from text prompts. We leverage panoramic videos as an intermediate representation to model the 360° details of a scene. WorldPrompter incorporates a conditional 360° panoramic video generator, capable of producing a 128-frame video that simulates a person walking through and capturing a virtual environment. The resulting video is then reconstructed as Gaussian splats by a fast feedforward 3D reconstructor, enabling a true walkable experience within the 3D scene. Experiments demonstrate that our panoramic video generation model achieves convincing view consistency across frames, enabling high-quality panoramic Gaussian splat reconstruction and facilitating traversal over an area of the scene. Qualitative and quantitative results also show it outperforms the state-of-the-art 360° video generators and 3D scene generation models.
Figure 1: We present WorldPrompter, a method that generates traversable environments from a text prompt alone (top). Our method first generates a spatially coherent video (2nd row) that travels through the environment described by the prompt. Next, we elevate the scene into a 3D Gaussian Splat (bottom), from which we project some views into standard perspective images (3rd row).
*Work done as an intern at Adobe Research.
1 Introduction
Scene-level 3D generation has attracted increasing research interest in recent years. For example, methods like RealmDreamer [38] and LucidDreamer [10] can generate Gaussian Splatting (GS) 3D scenes from text prompts. They rely on single-view diffusion models and monocular depth estimation to generate scenes progressively from one view to another by 2D inpainting. However, the 3D scenes created by these methods are confined to a limited viewable area and are not freely traversable. This constraint differentiates these approaches from a ”complete” scene generation that offers a navigable virtual world, similar to what is found in interactive video games.
In this paper, we propose WorldPrompter, the first generative pipeline that produces a navigable 3D scene from text prompts. Training a model to directly synthesize a full scene can be very challenging due to highly scarce scene-level 3D data. Drawing inspiration from conventional 3D capture methods and recent advancements in video generation models ([32, 54, 4]), we propose using videos as an intermediate representation to model complete 3D environments. Specifically, we train a diffusion model to generate videos traversing the environment based on text prompts. Moreover, instead of synthesizing captures from perspective cameras, we leverage panoramic cameras that directly capture the entire surroundings of an environment, providing 360° viewpoints that are ideal for our application.
Specifically, WorldPrompter is a two-stage pipeline composed of a 360° panoramic video generation model and a 3D reconstruction step to generate the final 3D Gaussians, a summary of which is shown in Fig. 1. The text-conditioned panoramic video generation model is trained to generate a 360° panoramic video that simulates a person walking through an environment while holding a panoramic camera to capture the scene. To train the video generator, we curated a dataset comprising approximately 1,700 panoramic videos and a panorama image dataset containing more than 200,000 high-quality panoramic in-the-wild captures. To enhance the generative capability of our method, we fine-tuned a Diffusion Transformer (DiT)-based video generator on our dataset to produce 10.6-second 360° panoramic videos (128 frames at 12 fps). We introduce a masked diffusion loss to ensure a clean panoramic capture sequence free from the person holding the camera, ready for subsequent reconstruction purposes. From the generated videos, we extract panoramic frames and project them from their original spherical coordinates onto the image plane to obtain perspective images from multiple viewpoints throughout the scene. The reconstruction step then builds the final 3D Gaussian Splat (3DGS) scene from these cropped images using a fast feedforward reconstruction model [56]. The resulting scene allows users to navigate freely within a virtual environment.
We show that our 360° video generation model achieves state-of-the-art generation capability in view coverage, temporal consistency, and visual quality. The high view consistency enables the videos to be reliably used as proxies for 3D scene generation, resulting in a significantly better view coverage of the generated scene and a traversable scene, more flexible than prior art. We highlight our contributions as follows:
  • We present WorldPrompter, the first generative pipeline to synthesize a complete 3D scene from a text prompt.
  • We propose leveraging panoramic capture videos as intermediate proxies for full 3D scene generation, addressing the scarcity of complete 3D scene datasets.
  • We introduce a high-quality 360° panoramic video generation model trained with mixed datasets and a masked diffusion loss achieving superior temporal consistency and view coverage compared to previous work.
  • We show that our 360° video is consistent enough to be efficiently turned into a 3DGS scene.
Figure 2: Overview of WorldPrompter. (top) We train a text-to-video model on a mix of 360° videos and images depicting in-the-wild environments. As it is challenging to avoid the person and camera equipment being visible in the video capture, we mask these elements out of the frame using a pretrained image segmentation model [1], and obtain the prompts from the video frames using LLaVA [24] for the videos, and BLIP-2 [19] for images. (bottom) At inference time, a user supplies a text prompt to our text-to-video model, which produces a “walk-through” video of the scene, which we reconstruct into a 3D Gaussian splat representation using Long-LRM [57].
2 Related Work
2.1 Panoramic Image and Video Generation
In computer graphics, 360° panoramas—often called IBLs or environment maps—are widely used to represent distant lighting and environments [12]. Numerous studies estimate lighting as High Dynamic Range (HDR) panoramas from input photos [51, 44, 11]. However, the resolution of these panoramas is often limited, making them suitable for light representation but insufficient for 3D reconstruction.
Advancements in diffusion models have made high-resolution image synthesis possible [35, 13]. Panorama generation also benefits from the use of diffusion models. For example, Text2Light [9] generates 4K HDR panoramas via text-driven low-resolution generation and super-resolution inverse tone mapping. MVDiffusion [39] creates panoramas by stitching consistent multi-view images generated from text prompts using attention layers. PanFusion [52] employs a dual-branch diffusion model with equirectangular-Perspective Projection Attention (EPPA) to improve consistency and control across panorama and perspective domains.
Diffusion models have also demonstrated their potential in video generation [6, 5]. Diffusion Transformers [31] offer a more versatile method for modeling high-dimensional data, making them particularly well-suited for video generation [7]. For 360° video generation, Wang et al. propose 360DVD for generating 360° panoramic videos from text prompts and motion conditions, where they leverage a lightweight 360-Adapter to finetune the Stable Diffusion [35] text-to-image model. However, their model is limited to generating only 16 frames, whereas our 360° video generation model produces significantly more frames with superior temporal consistency.
2.2 3D Neural Representation and Reconstruction
3D reconstruction and representation is an extensively studied field. While mesh-based methods [17] and point-based methods [33] have been extensively explored, they are limited in the quality of their results. In contrast, neural representations offer a flexible way to reconstruct 3D data. For instance, Neural Radiance Fields (NeRF) [26] introduces a novel way to represent 3D scenes by encoding volumetric scene information as a continuous function parameterized by a neural network, which inspired many follow-up works [27, 2, 3] that optimize efficiency, quality, and rendering speed. However, the MLP-based paradigm requires training a distinct model per scene, making reconstruction time-consuming. Additionally, rendering views involves model queries, creating a bottleneck for real-time applications.
3D Gaussian Splatting (3DGS) [18] introduces a memory-efficient and low-overhead 3D implicit representation using Gaussians derived from Structure-from-Motion (SfM) points. By combining anisotropic covariance optimization with a fast visibility-aware rendering algorithm, 3DGS enables efficient scene representation, high-quality real-time rendering, and has quickly emerged as a new paradigm in neural reconstruction [49, 47, 16, 14, 8]. Nevertheless, 3DGS still requires an optimization process, which can be costly for larger scenes. Recent advances such as Large Reconstruction Models (LRMs) [56, 45, 48, 53, 41] target scene reconstruction with a single feed-forward evaluation during inference. They achieve a reconstruction quality comparable to optimization-based methods while significantly improving their efficiency. In our scene generation pipeline, we also utilize this state-of-the-art feed-forward reconstruction approach for long-sequence reconstruction, achieving fast reconstruction.
2.3 3D Scene Generation
3D generation has emerged as a prominent focus in generative modeling. Given the scarcity of high-quality 3D data, recent works usually leverage intermediate representations, such as multi-view images, to improve the quality and efficiency of their reconstructions. Many studies have explored generating 3D objects in various formats, including Gaussian splats and meshes. For instance, Instant3D [20] accelerates 3D asset creation from text prompts by combining sparse-view generation with transformer-based reconstruction. Similarly, MVDream [37] improves consistency through multi-view diffusion by integrating 2D and 3D data. One-2-3-45 [25] simplifies single-image 3D reconstruction, enabling view-consistent mesh generation.
Scene-level generation, however, is a more challenging task because it requires more observations to achieve comprehensive view coverage. Typically, 3D scene generation methods employ progressive pipelines [29, 10, 15]: anchor cameras are first selected to generate text-aligned 2D images and corresponding depth maps, followed by a neural reconstruction process to build the initial 3D structure. This process is iteratively repeated to expand scene coverage by generating additional camera views and inpainting missing regions. However, such methods often suffer from prolonged generation times and incomplete coverage. Other studies have explored interactive and layout-guided approaches for 3D scene generation. Systems like WonderWorld [50] provide user-friendly interfaces for scene customization, while Director3D [22] facilitates real-world scene creation with adaptive camera trajectories.
The closest method to ours is DreamScene360 [55], which generates 3D scenes from text prompts using a single panorama image, along with depth estimation and Gaussian Splats optimization. However, while the resulting scene supports 360° viewing, it suffers from a limited movable range. 4K4DGen [21], introduces panorama-to-4D generation but still remains limited in terms of traversable area. More recently, ODIN [40] proposed learning 3D from 360° videos. However, their approach extracts only cropped views from videos to perform image-to-image style novel view synthesis, under-utilizing the full potential of 360° videos. As a result, ODIN fails to model a complete 3D world and can only generate partial scenes. In contrast to these approaches, our model generates complete scenes by fully leveraging the 360° representation, ensuring consistency and comprehensive coverage through a panoramic video generation model.
3 WorldPrompter
3.1 Overview
Our WorldPrompter synthesizes a traversable, 360° viewable 3D scene in two stages: generation and reconstruction. The generation stage leverages a text-conditioned 360° panoramic video diffusion model to generate a 10.6-second panoramic capture video for a target scene, simulating someone holding the camera and walking through the environment. Later, the reconstruction stage runs COLMAP [36] to estimate precise camera poses from perspective crops of the generated videos and builds a navigable 3DGS using Long-LRM [56] with the calibrated camera poses. The generated panoramic video features rich details and geometrically consistent visual details from multiple viewpoints along the traversal trajectory, ensuring high-quality reconstruction by the reconstruction step and yielding appealing novel views in the traversable scene.
3.2 Video Generator for 360° Panoramic Captures
Training a generative model for 360° videos is challenging due to limited data; therefore, we start with an existing video generator. We leverage a pretrained text-to-video diffusion model (similar to [7]) as backbone, and fine-tune it using panoramic data.
Specifically, our model’s backbone architecture $f_{\theta}$ is based on Diffusion Transformers [31]. Typically, in DiT models, a video sequence $x\in\mathbb{R}^{T\times 3\times H\times W}$ is first tokenized into patches, yielding a set of visual tokens $\hat{x}\in\mathbb{R}^{K\times D}$, where $T$, $H$, and $W$ correspond to the temporal and spatial dimensions of the video, $K$ represents the total number of tokens, and $D$ denotes the feature dimension. Being a text-conditioned model, it requires text prompts $c$ that we encode using a pre-trained model [34] and map into the same feature space $D$ by an embedding layer. The transformer $f_{\theta}(\hat{x}_{t};c,t)$ operates on these inputs, denoising the tokens at each timestep $t$. The denoised tokens $\hat{x}_{0}\in\mathbb{R}^{K\times D}$ are subsequently decoded and reassembled into the video data $x_{0}\in\mathbb{R}^{T\times 3\times H\times W}$ via linear layers. To reduce the computational complexity caused by the high-dimensionality of video data, the model performs its denoising in latent space after encoding by a Variational AutoEncoder (VAE) [31, 35].
The scarcity of 360° video data presents a significant challenge in training this model. To develop our high-quality 360° video generator, we curate two carefully selected datasets: (1) a 360° video dataset capturing static indoor environments and (2) a 360° image dataset featuring diverse real-world scenes.
3.2.1 360° video dataset
We collect a 360° video dataset with approximately 1,700 panoramic videos. Each video is captured by a person holding a Ricoh Theta Z1 camera walking inside a nearly static indoor scene to ensure the quality of 3D scene reconstruction. Each video has around 50 seconds of capture, yielding roughly 1,500 frames. To build a dataset for training, we segment the longer video sequences into smaller chunks, typically a 128-frame video clip (5.3/10.6 second at 24/12fps). We use LLaVA [23] to generate detailed captions describing each scene.
3.2.2 360° image dataset
The 360° video data is primarily captured indoors to ensure a static environment with no moving objects during the capture process. However, this may bias the fine-tuned model, limiting its ability to generalize effectively to outdoor environments. To mitigate this issue, we gather another dataset with around 200,000 panoramic images taken with cameras on a tripod, without the photographer in the scene. We caption each image with Blip-2 [19] to describe the whole scene in detail. These stationary images, captured in-the-wild, represent a much broader variety of scenes, thereby enhancing the generalization capability of our video diffusion model.
3.2.3 Mixed Training with Masked Loss
We combine the two described datasets to train our model, enabling the text-to-360° video generator to produce both 360° videos and images, with images treated as single-frame videos. While panoramic images depict a clean environment without the photographer present, the 360° video capture often includes the photographer visible in the scene. This causes the trained model to produce camera operators in its outputs (see Fig. 7), complicating subsequent reconstruction steps. Hence, when training our diffusion models on video data, we introduce a masked diffusion loss, where the photographer is excluded from the loss computation.
To achieve this, we preprocess all 360° videos by masking out the photographer using a pretrained image segmentation model [1] on each frame in a video. For a 128-frame chunk, we merge all segmentation masks to minimize segmentation errors and ensure a clean region without the person. Additionally, we conservatively mask out the bottom region of the equirectangular panoramic frames to eliminate visible camera equipment and the photographer’s hand. Assuming the corresponding binary mask $M$ is resized into the latent space by nearest neighbor interpolation, the masked diffusion loss is computed as
$\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,1)}\left[\left\lVert M\odot\left(\epsilon_{t}-f_{\theta}(\hat{x}_{t};c,t)\right)\right\rVert^{2}\right],$ (1)
where $\epsilon_{t}$ is the sampled noise at timestep $t$, and $M$ has the same dimensions as $\epsilon_{t}$, where $M[i,j]=1$ includes the region in the loss calculation, and $M[i,j]=0$ excludes it (indicating the visible person).
Our video generator is finally trained on a mix of $1/3$ image data and $2/3$ video data, achieving an empirically reasonable balance. The video diffusion model generates a 128-frame video clip at 12 fps, equivalent to a 10.6-second 360° capture video at a native resolution of 352$\times$704 pixels. We further apply a pretrained GAN-based video upsampler [43] to super-resolve our generated videos 4 times, resulting a final 360° video of 1416x2832 resolution.
3.3 Reconstruction from “Generated” Captures
Given the generated 360° capture video with 128 panoramic frames, we reconstruct a 3D Gaussian scene. For each panoramic frame, we randomly perform 3 perspective crops with a field of view of 120° and resolution of 512$\times$512, resulting in 384 perspective images for pose estimation and reconstruction. Similar to standard scene reconstruction, we calibrate the camera poses of these perspective images using COLMAP [36]. The high average matching ratio $94.6\%$ from COLMAP estimation demonstrates the strong view consistency of the generated videos from our text-to-360° video generator. We adopt Long-LRM, the state-of-the-art feed-forward 3DGS reconstructor [56] to reconstruct the final 3D scene using the estimated camera poses and generated views. Using Long-LRM allows us to achieve fast reconstruction compared to an optimization-based approach. We randomly subsample 32 random views along the original camera path as the inputs since the Long-LRM has a maximum limit for input views. The calibrated camera poses from COLMAP are converted into Plücker rays as the inputs for long-LRM.
3.4 Implementation Details
We fine-tune a pretrained text-to-video diffusion model using mixed 360° image and video datasets. The fine-tuning is conducted on 32 NVIDIA A100 GPUs. 8 GPUs train on 360° images, while 24 generate 128-frame 360° videos. Of these, 12 GPUs produce videos at 24 fps, and 12 generate videos at 12 fps. The model was finetuned from a pretrained text-to-video checkpoint for 20,000 steps in around 200 hours. During inference, we generate videos at 12 fps to extend the traversal range while maintaining good temporal consistency comparable to 24 fps generation.
The inference of a 10.6 second video takes around 10 minutes with an unoptimized PyTorch runtime [30] using 50 diffusion steps, together with video upsampling process. The entire reconstruction takes around 5 minutes, including pose estimation and Long-LRM reconstruction. Our full 3D generation pipeline requires approximately 15 minutes to generate a detailed GS scene, with potential for further optimization to improve inference speed and efficiency for different components.
[…]
360DVD [42] 0.7643 3.5797 NA
DreamScene360 [55] 0.8116 4.0942 2.4125
Ours 0.7387 4.2444 2.8043
Table 1: We quantitatively evaluate our method against the state-of-the-art (SOTA) 360° video generation model [42] and the SOTA text-to-scene model [55], which generates 360° coverage scenes. CLIP distance and Q-Align metrics are computed on the generated 360° images or frames across 320 test text prompts. Additionally, Q-Align NVS assesses the quality of novel view synthesis.
A calm Zen rock garden in an ancient monastery courtyard.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A candlelit medieval cathedral crypt with carved statues.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A Celtic stone circle on rolling green hills at dawn.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A glacier-capped polar landscape with drifting icebergs.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A cramped submarine cafeteria with portholes showing ocean depths.
Refer to caption (a) Panorama Reference Refer to caption (b) Novel View 1 Refer to caption (c) Novel View 2 Refer to caption (d) Novel View 3 Refer to caption (e) Novel View 4
Figure 3: Scene generation results. We present our scene generation results by showing (1) the first 360° image from our generated panoramic video as reference, and (2) novel video renderings (Novel View 1-4) from different spatial locations and viewpoints, demonstrating that our approach effectively synthesizes a traversable, 360° viewable scene. More rendered videos and navigable Gaussian Splats scenes can be found in the supplementary documents.
A cottage in the woods where the walls are made of interwoven tree roots, and the floor is a carpet of moss.
DreamScene360
[55]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
A cottage surrounded by fields of wildflowers that change color with the seasons, glowing faintly at twilight.
DreamScene360
[55]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
(a) Panorama Reference
(b) Novel View 1
(c) Novel View 2
(d) Novel View 3
(e) Novel View 4
Figure 4: Comparison of scene generation results. We compare our method with DreamScene360 [55] for text-based scene generation, as both approaches use panoramas as an intermediate representation for generating 360° viewable scenes. The first column presents the panorama reference: for DreamScene360, it is a single generated panorama from a text-to-image model, whereas in our results, it is the first frame of our generated 360° panorama video. While the previous work can produce a 360° viewable scene, its range of motion is highly limited. When novel viewpoints (Novel View 1-4) deviate significantly from the initial position, their GS scene exhibits severe blurriness and artifacts. Besides, our method achieves better alignment with the input prompt and produces the correct global structure of the scene.
4 Results
We demonstrate that our pipeline achieves state-of-the-art quality in both 360° video generation (Sec. 4.1) and 3D scene generation (Sec. 4.2), offering significantly improved visual consistency and view coverage. More visual examples, including videos and Gaussian Splats with an interactive viewer, can be found in the supplementary material. To quantitatively evaluate our method, we prompt ChatGPT [28] to generate a list of 320 descriptions of 3D scenes. We then use CLIP distance [34] to assess prompt adherence and Q-Align [46] to evaluate visual quality.
Additionally, we perform ablation studies in Sec. 4.3 to evaluate our design choices, including mixed data training and masked loss, to ensure the generation of high-quality videos suitable for reconstruction.
4.1 Text-to-360°-Video Generation
Our generation model can generate temporally consistent 360° panoramic videos, exceeding the quality of the state-of-the-art panoramic video generation models. We compare our 360° video generation results with 360DVD [42] in Fig. 5. We showcase four evenly sampled frames comparing 360DVD and our 360° panorama video generation model. Our model generates videos that closely match the input text prompt and produce richer scene details compared to 360DVD. The higher average CLIP distance and Q-Align metric for scenes generated from the 320 prompts in Table 1 further demonstrate the superior prompt alignment and visual quality of our method.
Additionally, videos generated by our model cover a broader scene span compared to 360DVD, which exhibits only slight rotations and translations of the viewpoint as indicated in the marked green box (Fig. 5). More importantly, since the primary goal of 360° panoramic video synthesis is to generate a coherent 3D scene, the quality of 360DVD’s results prevent it from effectively accomplishing this task. Its outputs lack structural consistency across views, exhibiting continuous morphing and noticeable changes, which undermine the perception of a stable 3D environment.
To validate this statement, we run COLMAP on their generated panoramas using a similar cropping and pose estimation procedure as described in Sec. 3.3. We find that the failure rate of COLMAP reaches $89.7\%$, and even in cases where COLMAP succeeds, the match rate remains very low, with only $41.8\%$ images being successfully matched, making their results unsuitable for 3D reconstruction. In contrast, our model preserves view-consistent local structures across frames. We do not encounter any failure cases when running COLMAP, and the average image match rate is as high as $94.6\%$ across our test prompts. This is the key reason we achieve a well-reconstructed 3D scene.
Ours DreamScene360 [55] LucidDreamer [10]
View Coverage 360° 360° Wide-angle front view
Movable Range Large Very Limited Limited
Table 2: Our method achieves state-of-the-art view coverage and traversable range compared to previous 3D generation methods.
4.2 Text-to-Scene Generation
As shown in Fig. 3, by chaining our 360° video generator with the 3D reconstruction process, we enable the creation of a diverse set of 3D scenes, allowing users to navigate the environment from novel viewpoints and spatial locations. More results can be found in the supplementary material.
In Fig. 4, we compare our results to the previous state-of-the-art text-to-scene method that can also create 360° virtual environment. However, as an approach that reconstructs a scene from only a single panorama, DreamScene360 cannot create a truly navigable scene. When synthesizing novel views (Novel View 1-4 in Fig. 4) from locations different from the initial position, the rendered images exhibit significant artifacts and blurriness. In Table 1, we compute the Q-align metrics on the rendered novel views (Q-align NVS), where our results achieve higher visual quality and less blurriness compared to previous work, demonstrating that our generated scene enables a more realistic navigation experience. Similarly, we also report the CLIP distance and Q-Align metrics on the generated panorama images, demonstrating that our method generates a 3D environment that more accurately responds to user requests while achieving superior visual quality, even at the initial position.
In addition to quantitative measurements, we highlight that our method is the first to achieve traversable and 360° viewable text-to-scene generation. As shown in Table 2, compared to previous work, our method achieves a complete 360° viewing area while also allowing users to walk inside the generated 3D scene for a large range.
4.3 Ablation Study
Training a high-quality 360° video generator within our 3D generative pipeline is a key factor in the success of our method, relying on the use of mixed training data and the masked diffusion loss. We conduct ablation studies on these design choices to demonstrate their effectiveness.
4.3.1 Mixed Training Data
We conduct an experiment by excluding the additional image panorama data. As shown in Fig. 6, incorporating image data significantly enhances the generalization capability of the video generator across a wide variety of scenes, and improves visual quality. Notably, due to the masked diffusion loss training, ground-truth pixels in the bottom regions of the panoramic videos are unavailable. The image data compensate for this limitation, enabling the model to intelligently auto-inpaint the bottom regions, resulting in a natural and realistic appearance.
4.3.2 Masked Diffusion Loss
We also train a model using the full diffusion loss applied to all pixels. The generated videos in this case reveal the photographer and the camera pod, which consistently appear in the 360° video training data (Fig. 7). This obstructs a significant portion of the generated capture videos, making comprehensive scene reconstruction challenging. By applying the masked diffusion loss and incorporating the aforementioned image panorama data, we guide the model to generate clean 360° videos that can be directly used for perspective cropping and reconstruction.
A cozy cabin library overlooking a vast snowy canyon.
360DVD [42] Refer to caption Refer to caption Refer to caption Refer to caption
Ours
A coastal castle perched on a cliff under a dramatic sunset.
360DVD [42] Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Figure 5: Comparison of 360° video generation results. We present a comparison of 360° panoramic video generation results against 360DVD [42]. Our generated videos exhibit significantly better visual quality and prompt alignment. We manually mark an anchor region with a green box in the 360° image, highlighting the substantially longer traversal range of our generated videos.
A cozy cottage surrounded by wildflowers that glow in the moonlight, nestled deep in an enchanted forest.
w/o Image Data Refer to caption Refer to caption
w/ Image Data
A serene desert where the sand shifts in swirling patterns, reflecting the colors of the sunset.
w/o Image Data Refer to caption Refer to caption
w/ Image Data
Figure 6: We verify that incorporating additional panorama image data helps the model generalize to diverse environments. Without the image data, the generated 360° videos not only exhibit degraded visual appearance but also display messy and incorrect content at the bottom of the video. This issue arises because these pixels lack supervision due to the masked loss.
A garden where every flower blooms in perfect circles, their petals reflecting the colors of the stars above.
Baseline Refer to caption Refer to caption
w/ Masked Loss
A room where the ceiling is made of glass, revealing the swirling, ever-changing night sky beyond.
Baseline Refer to caption Refer to caption
w/ Masked Loss
Figure 7: We conduct an ablation study to evaluate the impact of the masked diffusion loss. This loss function effectively guides the model to generate only the pixels of the scene, eliminating the moving photographer present in the original video training data. As a result, it enables the generation of a complete 360° scene capture for reconstruction.
5 Conclusion
In conclusion, we presented WorldPrompter, the first generative pipeline capable of producing a navigable 3D scenes from text prompts. By leveraging 360° panoramic video generation as an intermediate representation and a fast 3D reconstruction model, WorldPrompter enables the synthesis of immersive environments with comprehensive view coverage and high visual quality. This approach bridges the gap between text-to-3D generation and full scene synthesis.
Although WorldPrompter achieves state-of-the-art performance, generating longer panoramic videos, spanning minutes rather than seconds, remains an open challenge. Consistent content creation over extended sequences is crucial for scaling up to larger, more detailed 3D scenes. Addressing this challenge would further expand applications in virtual worlds and beyond, paving the way for future advancements in scene-level generative models.

### v4

@@
-WorldPrompter: Traversable Text-to-Scene Generation
-Abstract
-Scene-level 3D generation is a challenging research topic, with most existing methods generating only partial scenes and offering limited navigational freedom.
-We introduce WorldPrompter, a novel generative pipeline for synthesizing traversable 3D scenes from text prompts.
-We leverage panoramic videos as an intermediate representation to model the 360° details of a scene.
+Generating 360° Video is What You Need For a 3D Scene
+Abstract.
+Generating 3D scenes is still a challenging task due to the lack of readily available scene data.
+Most existing methods only produce partial scenes and provide limited navigational freedom.
+We introduce a practical and scalable solution that uses 360° video as an intermediate scene representation, capturing the full-scene context and ensuring consistent visual content throughout the generation.
+We propose WorldPrompter, a generative pipeline that synthesizes traversable 3D scenes from text prompts.
 WorldPrompter incorporates a conditional 360° panoramic video generator, capable of producing a 128-frame video that simulates a person walking through and capturing a virtual environment.
 The resulting video is then reconstructed as Gaussian splats by a fast feedforward 3D reconstructor, enabling a true walkable experience within the 3D scene.
-Experiments demonstrate that our panoramic video generation model achieves convincing view consistency across frames, enabling high-quality panoramic Gaussian splat reconstruction and facilitating traversal over an area of the scene.
+Experiments demonstrate that our panoramic video generation model, trained with a mix of image and video data, achieves convincing spatial and temporal consistency for static scenes.
+This is validated by an average COLMAP matching rate of 94.6%, allowing for high-quality panoramic Gaussian splat reconstruction and improved navigation throughout the scene.
 Qualitative and quantitative results also show it outperforms the state-of-the-art 360° video generators and 3D scene generation models.
-Figure 1: We present WorldPrompter, a method that generates traversable environments from a text prompt alone (top).
-Our method first generates a spatially coherent video (2nd row) that travels through the environment described by the prompt.
-Next, we elevate the scene into a 3D Gaussian Splat (bottom), from which we project some views into standard perspective images (3rd row).
-*Work done as an intern at Adobe Research.
-1 Introduction
+Diffusion Models, 360° Video Generative Model, 3D Reconstruction
+submissionid: 1445journal: TOG
+Figure 1.
+We present WorldPrompter, a method that generates traversable environments from a text prompt alone (above).
+Our method first generates a spatially coherent 360° video capture (top) that travels through a static environment described by the prompt.
+Next, we elevate generated video capture into a 3D Gaussian Splat (middle), enabling free navigation within a 3D scene using a virtual perspective camera to generate novel views (bottom).
+1. Introduction
 Scene-level 3D generation has attracted increasing research interest in recent years.
-For example, methods like RealmDreamer [38] and LucidDreamer [10] can generate Gaussian Splatting (GS) 3D scenes from text prompts.
+The limited availability of data generally hinders direct 3D scene generation (Liu et al., 2024a), making it difficult for methods to generalize to realistic scene synthesis effectively.
+To address this challenge, recent approaches typically rely on an image diffusion model to create realistic scene views.
+For example, methods like RealmDreamer (Shriram et al., 2025) and LucidDreamer (Chung et al., 2023) can generate Gaussian Splatting (GS) 3D scenes from text prompts.
 They rely on single-view diffusion models and monocular depth estimation to generate scenes progressively from one view to another by 2D inpainting.
@@
 This constraint differentiates these approaches from a ”complete” scene generation that offers a navigable virtual world, similar to what is found in interactive video games.
-In this paper, we propose WorldPrompter, the first generative pipeline that produces a navigable 3D scene from text prompts.
-Training a model to directly synthesize a full scene can be very challenging due to highly scarce scene-level 3D data.
-Drawing inspiration from conventional 3D capture methods and recent advancements in video generation models ([32, 54, 4]), we propose using videos as an intermediate representation to model complete 3D environments.
+Recent works (Zhou et al., 2024; Yang et al., 2024) attempt to bridge the gap between partial and full scene generation by leveraging panoramic images.
+However, these methods still rely on outpainting to fill in missing regions, which inherently limits the navigable range and often leads to distortions during viewpoint transitions.
+More importantly, we observe that scenes generated via these outpainting-based approaches frequently suffer from distortion and inconsistencies in visual content across the scene.
+We hypothesize that these issues stem from the commonly used progressive outpainting approach, which independently synthesizes new content based on the previously generated content.
+However, this assumption is flawed, as visual content is globally interconnected and consistent rather than strictly sequential (Tian et al., 2024).
+As a result, while their generated scenes may appear locally plausible, they exhibit significant global distortions (Fig.
+7).
+In this paper, instead of synthesizing a scene in a view-by-view fashion, we aim to generate a maximally navigable scene in a single pass, ensuring undistorted global visual content synthesis.
+Although scenes can be represented in various formats, such as meshes or point clouds, data in these formats is often limited and difficult to scale.
+Drawing inspiration from conventional 3D capture methods and recent advancements in video generation models ((Polyak et al., 2024; Zheng et al., 2024; Blattmann et al., 2023a)), we propose using videos as an intermediate representation to model complete 3D environments.
+Modeling scenes using video captures provides a scalable approach for creating training data, enabling generative models to achieve strong generalization capabilities.
+To validate this idea, we propose WorldPrompter, a generative pipeline that produces a navigable 3D scene from text prompts.
 Specifically, we train a diffusion model to generate videos traversing the environment based on text prompts.
 Moreover, instead of synthesizing captures from perspective cameras, we leverage panoramic cameras that directly capture the entire surroundings of an environment, providing 360° viewpoints that are ideal for our application.
-Specifically, WorldPrompter is a two-stage pipeline composed of a 360° panoramic video generation model and a 3D reconstruction step to generate the final 3D Gaussians, a summary of which is shown in Fig.
+WorldPrompter is a two-stage pipeline composed of a 360° panoramic video generation model and a 3D reconstruction step to generate the final 3D Gaussians, a summary of which is shown in Fig.
 1.
 The text-conditioned panoramic video generation model is trained to generate a 360° panoramic video that simulates a person walking through an environment while holding a panoramic camera to capture the scene.
-To train the video generator, we curated a dataset comprising approximately 1,700 panoramic videos and a panorama image dataset containing more than 200,000 high-quality panoramic in-the-wild captures.
+To train the video generator, we collected a dataset comprising approximately 1,700 panoramic videos and a panorama image dataset containing more than 200,000 high-quality panoramic in-the-wild captures.
 To enhance the generative capability of our method, we fine-tuned a Diffusion Transformer (DiT)-based video generator on our dataset to produce 10.6-second 360° panoramic videos (128 frames at 12 fps).
@@
 From the generated videos, we extract panoramic frames and project them from their original spherical coordinates onto the image plane to obtain perspective images from multiple viewpoints throughout the scene.
-The reconstruction step then builds the final 3D Gaussian Splat (3DGS) scene from these cropped images using a fast feedforward reconstruction model [56].
-The resulting scene allows users to navigate freely within a virtual environment.
-We show that our 360° video generation model achieves state-of-the-art generation capability in view coverage, temporal consistency, and visual quality.
-The high view consistency enables the videos to be reliably used as proxies for 3D scene generation, resulting in a significantly better view coverage of the generated scene and a traversable scene, more flexible than prior art.
+The reconstruction step then builds the final 3D Gaussian Splat (3DGS) scene from these cropped images using a fast feedforward reconstruction model (Ziwen et al., 2024).
+The resulting scene allows users to navigate freely within a virtual environment, beyond the original video camera path.
+We demonstrate that our 360° video generation model—trained on a mixture of image and video data—achieves state-of-the-art performance in temporal consistency and visual quality.
+The high consistency of the views, confirmed by a matching rate of up to 94.5% using COLMAP (Schönberger and Frahm, 2016), ensures that the videos can be reliably used as proxies for 3D scene generation.
+This results in a traversable, 360° viewable environment with globally coherent visual content.
+We demonstrate that our generated scene allows longer navigation range and offers superior global visual coherence and quality compared to prior art.
 We highlight our contributions as follows:
-  • We present WorldPrompter, the first generative pipeline to synthesize a complete 3D scene from a text prompt.
-  • We propose leveraging panoramic capture videos as intermediate proxies for full 3D scene generation, addressing the scarcity of complete 3D scene datasets.
-  • We introduce a high-quality 360° panoramic video generation model trained with mixed datasets and a masked diffusion loss achieving superior temporal consistency and view coverage compared to previous work.
-  • We show that our 360° video is consistent enough to be efficiently turned into a 3DGS scene.
-Figure 2: Overview of WorldPrompter.
+  • We propose leveraging panoramic capture videos as an intermediate 3D scene representation.
+We found it is a practical solution for scaling up training data for high-quality photorealistic scene synthesis.
+  • We present WorldPrompter, a practical generative text-to-scene pipeline that synthesizes a traversable 3D scene with strong global visual coherence.
+  • We introduce a high-quality 360° panoramic video generation model, where the proposed mixed datasets and masked diffusion loss play a critical role in achieving artifact-free and generalizable static 360° scene generation.
+Figure 2.
+Overview of WorldPrompter.
 (top) We train a text-to-video model on a mix of 360° videos and images depicting in-the-wild environments.
-As it is challenging to avoid the person and camera equipment being visible in the video capture, we mask these elements out of the frame using a pretrained image segmentation model [1], and obtain the prompts from the video frames using LLaVA [24] for the videos, and BLIP-2 [19] for images.
-(bottom) At inference time, a user supplies a text prompt to our text-to-video model, which produces a “walk-through” video of the scene, which we reconstruct into a 3D Gaussian splat representation using Long-LRM [57].
-2 Related Work
-2.1 Panoramic Image and Video Generation
-In computer graphics, 360° panoramas—often called IBLs or environment maps—are widely used to represent distant lighting and environments [12].
-Numerous studies estimate lighting as High Dynamic Range (HDR) panoramas from input photos [51, 44, 11].
+As it is challenging to avoid the person and camera equipment being visible in the video capture, we mask these elements out of the frame using a pretrained image segmentation model (Adobe, 2025), and obtain the prompts from the video frames using LLaVA (Liu et al., 2023b) for the videos, and BLIP-2 (Li et al., 2023a) for images.
+(bottom) At inference time, a user supplies a text prompt to our text-to-video model, which produces a “walk-through” video of the scene, which we reconstruct into a 3D Gaussian splat representation using Long-LRM (Ziwen et al., 2024).
+2. Related Work
+2.1. Panoramic Image and Video Generation
+In computer graphics, 360° panoramas—often called IBLs or environment maps—are widely used to represent distant lighting and environments (Debevec and Malik, 2023).
+Numerous studies estimate lighting as High Dynamic Range (HDR) panoramas from input photos (Zhan et al., 2021; Weber et al., 2022; Dastjerdi et al., 2023).
 However, the resolution of these panoramas is often limited, making them suitable for light representation but insufficient for 3D reconstruction.
-Advancements in diffusion models have made high-resolution image synthesis possible [35, 13].
+Advancements in diffusion models have made high-resolution image synthesis possible (Rombach et al., 2022; Esser et al., 2024).
 Panorama generation also benefits from the use of diffusion models.
-For example, Text2Light [9] generates 4K HDR panoramas via text-driven low-resolution generation and super-resolution inverse tone mapping.
-MVDiffusion [39] creates panoramas by stitching consistent multi-view images generated from text prompts using attention layers.
-PanFusion [52] employs a dual-branch diffusion model with equirectangular-Perspective Projection Attention (EPPA) to improve consistency and control across panorama and perspective domains.
-Diffusion models have also demonstrated their potential in video generation [6, 5].
-Diffusion Transformers [31] offer a more versatile method for modeling high-dimensional data, making them particularly well-suited for video generation [7].
+For example, Text2Light (Chen et al., 2022) generates 4K HDR panoramas via text-driven low-resolution generation and super-resolution inverse tone mapping.
+MVDiffusion (Tang et al., 2023) creates panoramas by stitching consistent multi-view images generated from text prompts using attention layers.
+PanFusion (Zhang et al., 2024b) employs a dual-branch diffusion model with equirectangular-Perspective Projection Attention (EPPA) to improve consistency and control across panorama and perspective domains.
+DiffPano (Ye et al., 2024a) fine-tunes Stable Diffusion (Rombach et al., 2022) using LoRA (Hu et al., 2021) to generate panorama images, and with a Spherical Epipolar-Aware Attention Module, it enables multi-view panorama generation.
+Diffusion models have also demonstrated their potential in video generation (Blattmann et al., 2023b, a).
+Diffusion Transformers (Peebles and Xie, 2023) offer a more versatile method for modeling high-dimensional data, making them particularly well-suited for video generation (Brooks et al., 2024).
 For 360° video generation, Wang et al.
-propose 360DVD for generating 360° panoramic videos from text prompts and motion conditions, where they leverage a lightweight 360-Adapter to finetune the Stable Diffusion [35] text-to-image model.
+propose 360DVD for generating 360° panoramic videos from text prompts and motion conditions, where they leverage a lightweight 360-Adapter to finetune the Stable Diffusion (Rombach et al., 2022) text-to-image model.
 However, their model is limited to generating only 16 frames, whereas our 360° video generation model produces significantly more frames with superior temporal consistency.
-2.2 3D Neural Representation and Reconstruction
+2.2. 3D Neural Representation and Reconstruction
 3D reconstruction and representation is an extensively studied field.
-While mesh-based methods [17] and point-based methods [33] have been extensively explored, they are limited in the quality of their results.
+While mesh-based methods (Kazhdan et al., 2006) and point-based methods (Qi et al., 2017) have been extensively explored, they are limited in the quality of their results.
 In contrast, neural representations offer a flexible way to reconstruct 3D data.
-For instance, Neural Radiance Fields (NeRF) [26] introduces a novel way to represent 3D scenes by encoding volumetric scene information as a continuous function parameterized by a neural network, which inspired many follow-up works [27, 2, 3] that optimize efficiency, quality, and rendering speed.
+For instance, Neural Radiance Fields (NeRF) (Mildenhall et al., 2020) introduces a novel way to represent 3D scenes by encoding volumetric scene information as a continuous function parameterized by a neural network, which inspired many follow-up works (Müller et al., 2022; Barron et al., 2022, 2023) that optimize efficiency, quality, and rendering speed.
 However, the MLP-based paradigm requires training a distinct model per scene, making reconstruction time-consuming.
 Additionally, rendering views involves model queries, creating a bottleneck for real-time applications.
-3D Gaussian Splatting (3DGS) [18] introduces a memory-efficient and low-overhead 3D implicit representation using Gaussians derived from Structure-from-Motion (SfM) points.
-By combining anisotropic covariance optimization with a fast visibility-aware rendering algorithm, 3DGS enables efficient scene representation, high-quality real-time rendering, and has quickly emerged as a new paradigm in neural reconstruction [49, 47, 16, 14, 8].
+3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) introduces a memory-efficient and low-overhead 3D implicit representation using Gaussians derived from Structure-from-Motion (SfM) points.
+By combining anisotropic covariance optimization with a fast visibility-aware rendering algorithm, 3DGS enables efficient scene representation, high-quality real-time rendering, and has quickly emerged as a new paradigm in neural reconstruction (Ye et al., 2024b; Wu et al., 2024; Jiang et al., 2024; Gao et al., 2024; Charatan et al., 2024).
 Nevertheless, 3DGS still requires an optimization process, which can be costly for larger scenes.
-Recent advances such as Large Reconstruction Models (LRMs) [56, 45, 48, 53, 41] target scene reconstruction with a single feed-forward evaluation during inference.
+Recent advances such as Large Reconstruction Models (LRMs) (Ziwen et al., 2024; Wei et al., 2024; Xie et al., 2024; Zhang et al., 2024a; Wang et al., 2023) target scene reconstruction with a single feed-forward evaluation during inference.
 They achieve a reconstruction quality comparable to optimization-based methods while significantly improving their efficiency.
 In our scene generation pipeline, we also utilize this state-of-the-art feed-forward reconstruction approach for long-sequence reconstruction, achieving fast reconstruction.
-2.3 3D Scene Generation
+2.3. 3D Scene Generation
 3D generation has emerged as a prominent focus in generative modeling.
-Given the scarcity of high-quality 3D data, recent works usually leverage intermediate representations, such as multi-view images, to improve the quality and efficiency of their reconstructions.
+Given the scarcity of high-quality 3D data, recent works usually leverage intermediate representations, such as multi-view images (Liu et al., 2023c; Voleti et al., 2024; Hu et al., 2024; Li et al., 2023b; Shi et al., 2024; Liu et al., 2024b) and videos (Parthasarathy et al., 2024; Han et al., 2024; Chen et al., 2024), to improve the quality and efficiency of their reconstructions.
 Many studies have explored generating 3D objects in various formats, including Gaussian splats and meshes.
-For instance, Instant3D [20] accelerates 3D asset creation from text prompts by combining sparse-view generation with transformer-based reconstruction.
-Similarly, MVDream [37] improves consistency through multi-view diffusion by integrating 2D and 3D data.
-One-2-3-45 [25] simplifies single-image 3D reconstruction, enabling view-consistent mesh generation.
+For instance, Instant3D (Li et al., 2023b) accelerates 3D asset creation from text prompts by combining sparse-view generation with transformer-based reconstruction.
+Similarly, MVDream (Shi et al., 2024) improves consistency through multi-view diffusion by integrating 2D and 3D data.
+One-2-3-45 (Liu et al., 2024b) simplifies single-image 3D reconstruction, enabling view-consistent mesh generation.
 Scene-level generation, however, is a more challenging task because it requires more observations to achieve comprehensive view coverage.
-Typically, 3D scene generation methods employ progressive pipelines [29, 10, 15]: anchor cameras are first selected to generate text-aligned 2D images and corresponding depth maps, followed by a neural reconstruction process to build the initial 3D structure.
+Typically, 3D scene generation methods employ progressive pipelines (Ouyang et al., 2023; Chung et al., 2023; Höllein et al., 2023): anchor cameras are first selected to generate text-aligned 2D images and corresponding depth maps, followed by a neural reconstruction process to build the initial 3D structure.
 This process is iteratively repeated to expand scene coverage by generating additional camera views and inpainting missing regions.
@@
 Other studies have explored interactive and layout-guided approaches for 3D scene generation.
-Systems like WonderWorld [50] provide user-friendly interfaces for scene customization, while Director3D [22] facilitates real-world scene creation with adaptive camera trajectories.
-The closest method to ours is DreamScene360 [55], which generates 3D scenes from text prompts using a single panorama image, along with depth estimation and Gaussian Splats optimization.
+Systems like WonderWorld (Yu et al., 2024a) provide user-friendly interfaces for scene customization, while Director3D (Li et al., 2024a) facilitates real-world scene creation with adaptive camera trajectories.
+Another potential limitation of this family of methods, as discussed in Sec.
+1, is the lack of guaranteed global visual coherence, often resulting in distorted scene layouts and object shapes.
+Most recent literature has discovered the potential of video generation for consistent scene-level generation.
+VideoScene (Wang et al., 2025) leverages the temporal consistency in video diffusion models to generate novel views for 3D reconstruction by interpolating two input keyframes.
+Yu et al.
+and Wu et al.
+explore explicitly reconstructing a 3D point cloud as spatial memory for consistent video generation.
+The closest method to ours is DreamScene360 (Zhou et al., 2024), which generates 3D scenes from text prompts using a single panorama image, along with depth estimation and Gaussian Splats optimization.
 However, while the resulting scene supports 360° viewing, it suffers from a limited movable range.
-4K4DGen [21], introduces panorama-to-4D generation but still remains limited in terms of traversable area.
-More recently, ODIN [40] proposed learning 3D from 360° videos.
-However, their approach extracts only cropped views from videos to perform image-to-image style novel view synthesis, under-utilizing the full potential of 360° videos.
+LayerPano3D (Yang et al., 2024) offers slightly increased navigational range by generating layered panoramas, but the content in the additional layers may suffer from distortion.
+ODIN (Wallingford et al., 2024) proposed learning 3D from 360° videos.
+However, their approach extracts only cropped views from videos to perform image-to-image style novel view synthesis, underutilizing the full potential of 360° videos.
 As a result, ODIN fails to model a complete 3D world and can only generate partial scenes.
-In contrast to these approaches, our model generates complete scenes by fully leveraging the 360° representation, ensuring consistency and comprehensive coverage through a panoramic video generation model.
-3 WorldPrompter
-3.1 Overview
-Our WorldPrompter synthesizes a traversable, 360° viewable 3D scene in two stages: generation and reconstruction.
+Beyond 3D, 4K4DGen (Li et al., 2024b), introduces panorama-to-4D generation but still remains limited in terms of traversable area.
+In contrast to these approaches, our model introduces 360° videos as an explicit 3D scene proxy in the generation task, offering greater data scalability and ensuring strong visual coherence alongside comprehensive view coverage, ultimately enabling a realistic navigation experience.
+3. WorldPrompter
+3.1. Overview
+We show that 360° video capture serves as an effective, all-you-need proxy for 3D scene generation, as validated by the WorldPrompter pipeline.
+It synthesizes a traversable, 360° viewable 3D scene in two stages: generation and reconstruction.
 The generation stage leverages a text-conditioned 360° panoramic video diffusion model to generate a 10.6-second panoramic capture video for a target scene, simulating someone holding the camera and walking through the environment.
-Later, the reconstruction stage runs COLMAP [36] to estimate precise camera poses from perspective crops of the generated videos and builds a navigable 3DGS using Long-LRM [56] with the calibrated camera poses.
-The generated panoramic video features rich details and geometrically consistent visual details from multiple viewpoints along the traversal trajectory, ensuring high-quality reconstruction by the reconstruction step and yielding appealing novel views in the traversable scene.
-3.2 Video Generator for 360° Panoramic Captures
-Training a generative model for 360° videos is challenging due to limited data; therefore, we start with an existing video generator.
-We leverage a pretrained text-to-video diffusion model (similar to [7]) as backbone, and fine-tune it using panoramic data.
-Specifically, our model’s backbone architecture $f_{\theta}$ is based on Diffusion Transformers [31].
-Typically, in DiT models, a video sequence $x\in\mathbb{R}^{T\times 3\times H\times W}$ is first tokenized into patches, yielding a set of visual tokens $\hat{x}\in\mathbb{R}^{K\times D}$, where $T$, $H$, and $W$ correspond to the temporal and spatial dimensions of the video, $K$ represents the total number of tokens, and $D$ denotes the feature dimension.
-Being a text-conditioned model, it requires text prompts $c$ that we encode using a pre-trained model [34] and map into the same feature space $D$ by an embedding layer.
-The transformer $f_{\theta}(\hat{x}_{t};c,t)$ operates on these inputs, denoising the tokens at each timestep $t$.
-The denoised tokens $\hat{x}_{0}\in\mathbb{R}^{K\times D}$ are subsequently decoded and reassembled into the video data $x_{0}\in\mathbb{R}^{T\times 3\times H\times W}$ via linear layers.
-To reduce the computational complexity caused by the high-dimensionality of video data, the model performs its denoising in latent space after encoding by a Variational AutoEncoder (VAE) [31, 35].
-The scarcity of 360° video data presents a significant challenge in training this model.
-To develop our high-quality 360° video generator, we curate two carefully selected datasets: (1) a 360° video dataset capturing static indoor environments and (2) a 360° image dataset featuring diverse real-world scenes.
-3.2.1 360° video dataset
-We collect a 360° video dataset with approximately 1,700 panoramic videos.
+Later, the reconstruction stage runs COLMAP (Schönberger and Frahm, 2016) to estimate precise camera poses from perspective crops of the generated videos and builds a navigable 3DGS using Long-LRM (Ziwen et al., 2024) with the calibrated camera poses.
+The generated panoramic video features rich details and geometrically consistent visual details over the entire scene, ensuring high-quality reconstruction and yielding appealing novel views in the traversable scene.
+3.2. Video Generator for 360° Panoramic Captures
+Training a generative model for 360° videos from scratch is challenging; therefore, we start with an existing video generator.
+We leverage a pretrained text-to-video diffusion model as backbone, and fine-tune it using panoramic data.
+Specifically, our model’s backbone architecture $f_{\theta}$ is based on Diffusion Transformers (Peebles and Xie, 2023).
+Typically, in latent DiT models, a video or an image is first encoded by a 3DVAE (Yu et al., 2024b), resulting a latent visual sequence $x\in\mathbb{R}^{T\times 3\times H\times W}$ ($T=1$ if it’s an image).
+The visual sequence is tokenized into patches, yielding a set of visual tokens $\hat{x}\in\mathbb{R}^{K\times D}$, where $T$, $H$, and $W$ correspond to the temporal and spatial dimensions of the video, $K$ represents the total number of tokens, and $D$ denotes the feature dimension.
+Being a text-conditioned model, it requires text prompts $c$ that we encode using a pre-trained model (Radford et al., 2021) and map into the same feature space $D$ by an embedding layer.
+As a decoder-only transformer architecture, the embedded text tokens are concatenated with the noised visual tokens.
+The transformer $f_{\theta}(\hat{x}_{t};c,t)$ operates on these inputs, denoising the visual tokens at each timestep $t$.
+The denoised tokens $\hat{x}_{0}\in\mathbb{R}^{K\times D}$ are subsequently decoded and reassembled into the visual data $x_{0}\in\mathbb{R}^{T\times 3\times H\times W}$ via linear layers, and finally decoded as videos or images.
+To preserve the generalization ability of the finetuned t2v model across diverse prompts, we curate two carefully selected datasets: (1) a 360° video dataset capturing static indoor environments, and (2) a 360° image dataset featuring a wide range of real-world scenes.
+We also incorporate a masked diffusion loss, which is key to ensuring artifact-free generation.
+3.2.1. 360° video dataset
+We start by collecting a 360° video dataset with approximately 1,700 panoramic videos.
 Each video is captured by a person holding a Ricoh Theta Z1 camera walking inside a nearly static indoor scene to ensure the quality of 3D scene reconstruction.
 Each video has around 50 seconds of capture, yielding roughly 1,500 frames.
-To build a dataset for training, we segment the longer video sequences into smaller chunks, typically a 128-frame video clip (5.3/10.6 second at 24/12fps).
-We use LLaVA [23] to generate detailed captions describing each scene.
-3.2.2 360° image dataset
-The 360° video data is primarily captured indoors to ensure a static environment with no moving objects during the capture process.
-However, this may bias the fine-tuned model, limiting its ability to generalize effectively to outdoor environments.
+To build a dataset for training, we segment the longer video sequences into smaller chunks, typically a 128-frame video clip (5.3/10.6 seconds at 24/12fps).
+We use LLaVA (Liu et al., 2023a) to generate detailed captions describing each scene.
+3.2.2. 360° image dataset
+However, as the 360° video data is primarily captured indoors to ensure a static environment with no moving objects during the capture process, this may bias the fine-tuned model, limiting its ability to generalize effectively to outdoor environments.
 To mitigate this issue, we gather another dataset with around 200,000 panoramic images taken with cameras on a tripod, without the photographer in the scene.
-We caption each image with Blip-2 [19] to describe the whole scene in detail.
+We caption each image with Blip-2 (Li et al., 2023a) to describe the whole scene in detail.
 These stationary images, captured in-the-wild, represent a much broader variety of scenes, thereby enhancing the generalization capability of our video diffusion model.
-3.2.3 Mixed Training with Masked Loss
+3.2.3. Mixed Training with Masked Loss
 We combine the two described datasets to train our model, enabling the text-to-360° video generator to produce both 360° videos and images, with images treated as single-frame videos.
@@
 This causes the trained model to produce camera operators in its outputs (see Fig.
-7), complicating subsequent reconstruction steps.
+4), complicating subsequent reconstruction steps.
 Hence, when training our diffusion models on video data, we introduce a masked diffusion loss, where the photographer is excluded from the loss computation.
-To achieve this, we preprocess all 360° videos by masking out the photographer using a pretrained image segmentation model [1] on each frame in a video.
+To achieve this, we preprocess all 360° videos by masking out the photographer using a pretrained image segmentation model (Adobe, 2025) on each frame in a video.
 For a 128-frame chunk, we merge all segmentation masks to minimize segmentation errors and ensure a clean region without the person.
@@
 Assuming the corresponding binary mask $M$ is resized into the latent space by nearest neighbor interpolation, the masked diffusion loss is computed as
-$\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,1)}\left[\left\lVert M\odot\left(\epsilon_{t}-f_{\theta}(\hat{x}_{t};c,t)\right)\right\rVert^{2}\right],$ (1)
+(1) $\mathbb{E}_{x\sim p_{\text{data}},t\sim U(0,1)}\left[\left\lVert M\odot\left(\epsilon_{t}-f_{\theta}(\hat{x}_{t};c,t)\right)\right\rVert^{2}\right],$
 where $\epsilon_{t}$ is the sampled noise at timestep $t$, and $M$ has the same dimensions as $\epsilon_{t}$, where $M[i,j]=1$ includes the region in the loss calculation, and $M[i,j]=0$ excludes it (indicating the visible person).
-Our video generator is finally trained on a mix of $1/3$ image data and $2/3$ video data, achieving an empirically reasonable balance.
-The video diffusion model generates a 128-frame video clip at 12 fps, equivalent to a 10.6-second 360° capture video at a native resolution of 352$\times$704 pixels.
-We further apply a pretrained GAN-based video upsampler [43] to super-resolve our generated videos 4 times, resulting a final 360° video of 1416x2832 resolution.
-3.3 Reconstruction from “Generated” Captures
+[…]
+3.3. Reconstruction from “Generated” Captures
 Given the generated 360° capture video with 128 panoramic frames, we reconstruct a 3D Gaussian scene.
-For each panoramic frame, we randomly perform 3 perspective crops with a field of view of 120° and resolution of 512$\times$512, resulting in 384 perspective images for pose estimation and reconstruction.
-Similar to standard scene reconstruction, we calibrate the camera poses of these perspective images using COLMAP [36].
+For each panoramic frame, we randomly perform 3 perspective crops with a field of view of 120° and resolution of $\times512$, resulting in 384 perspective images for pose estimation and reconstruction.
+Similar to standard scene reconstruction, we calibrate the camera poses of these perspective images using COLMAP (Schönberger and Frahm, 2016).
 The high average matching ratio $94.6\%$ from COLMAP estimation demonstrates the strong view consistency of the generated videos from our text-to-360° video generator.
-We adopt Long-LRM, the state-of-the-art feed-forward 3DGS reconstructor [56] to reconstruct the final 3D scene using the estimated camera poses and generated views.
+We adopt Long-LRM, the state-of-the-art feed-forward 3DGS reconstructor (Ziwen et al., 2024) to reconstruct the final 3D scene using the estimated camera poses and generated views.
 Using Long-LRM allows us to achieve fast reconstruction compared to an optimization-based approach.
@@
 The calibrated camera poses from COLMAP are converted into Plücker rays as the inputs for long-LRM.
-3.4 Implementation Details
+3.4. Implementation Details
 We fine-tune a pretrained text-to-video diffusion model using mixed 360° image and video datasets.
-The fine-tuning is conducted on 32 NVIDIA A100 GPUs.
-8 GPUs train on 360° images, while 24 generate 128-frame 360° videos.
-Of these, 12 GPUs produce videos at 24 fps, and 12 generate videos at 12 fps.
+The fine-tuning is conducted on 32 NVIDIA A100 GPUs with a mixed fps training schedule (12fps and 24fps).
 The model was finetuned from a pretrained text-to-video checkpoint for 20,000 steps in around 200 hours.
 During inference, we generate videos at 12 fps to extend the traversal range while maintaining good temporal consistency comparable to 24 fps generation.
-The inference of a 10.6 second video takes around 10 minutes with an unoptimized PyTorch runtime [30] using 50 diffusion steps, together with video upsampling process.
+The inference of a 10.6-second video takes around 10 minutes with an unoptimized PyTorch runtime (Paszke et al., 2019) using 50 diffusion steps, together with the video upsampling process.
 The entire reconstruction takes around 5 minutes, including pose estimation and Long-LRM reconstruction.
 Our full 3D generation pipeline requires approximately 15 minutes to generate a detailed GS scene, with potential for further optimization to improve inference speed and efficiency for different components.
+Figure 3.
+Comparison of 360° video generation results.
+We present a comparison of 360° panoramic video generation results against 360DVD (Wang et al., 2024).
+Our generated videos exhibit significantly better visual quality and prompt alignment.
+We manually mark an anchor region with a green box in the 360° image, highlighting the substantially longer traversal range of our generated videos.
+Stronger parallax is important for accurate 3D reconstruction.
+Figure 4.
+We conduct an ablation study to evaluate the impact of the masked diffusion loss.
+This loss function effectively guides the model to generate only the pixels of the scene, eliminating the moving photographer present in the original video training data.
+As a result, it enables the generation of a complete 360° scene capture for reconstruction.
+4. Results
+By generating 360° video capture, we demonstrate that our pipeline achieves state-of-the-art quality in 3D scene generation (Sec.
+4.2), offering significantly improved view coverage, traversability and coherent global visual content.
+This is primarily due to a strong 360° video generator for scenes (Sec.
+4.1).
+To quantitatively evaluate our method, we prompt ChatGPT (OpenAI, 2023) to generate a list of 320 descriptions of 3D scenes.
+Using this in-the-wild test set, we conduct systematic comparisons against existing methods, demonstrating superior video and 3D scene generation quality through both visual examples and quantitative metrics.
+More visual examples can be found in the supplementary material.
+Additionally, we perform ablation studies in Sec.
+4.3 to evaluate our design choices, including mixed data training and masked loss, to ensure the generation of high-quality videos suitable for artifact-free 3D scene reconstruction.
+Table 1.
+We quantitatively evaluate our method against the state-of-the-art (SOTA) 360° video generation model (Wang et al., 2024).
+We report CLIP distance (CLIP) for prompt alignment quality, Q-Align score for image quality, VBench for aesthetic/imaging quality, and COLMAP Failure Rate (COLMAP-FR) along with Matching Rate (COLMAP-MR) for multi-view consistency.
+The metrics are computed on the generated 360° frames across 320 test text prompts.
+Quantitative metrics across multiple dimensions demonstrate that our generated 360° videos achieve high visual quality and can be reliably used for 3D scene representation and reconstruction.
 […]
-360DVD [42] 0.7643 3.5797 NA
-DreamScene360 [55] 0.8116 4.0942 2.4125
-Ours 0.7387 4.2444 2.8043
-Table 1: We quantitatively evaluate our method against the state-of-the-art (SOTA) 360° video generation model [42] and the SOTA text-to-scene model [55], which generates 360° coverage scenes.
-CLIP distance and Q-Align metrics are computed on the generated 360° images or frames across 320 test text prompts.
-Additionally, Q-Align NVS assesses the quality of novel view synthesis.
-A calm Zen rock garden in an ancient monastery courtyard.
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-A candlelit medieval cathedral crypt with carved statues.
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-A Celtic stone circle on rolling green hills at dawn.
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-A glacier-capped polar landscape with drifting icebergs.
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-A cramped submarine cafeteria with portholes showing ocean depths.
-Refer to caption (a) Panorama Reference Refer to caption (b) Novel View 1 Refer to caption (c) Novel View 2 Refer to caption (d) Novel View 3 Refer to caption (e) Novel View 4
-Figure 3: Scene generation results.
-We present our scene generation results by showing (1) the first 360° image from our generated panoramic video as reference, and (2) novel video renderings (Novel View 1-4) from different spatial locations and viewpoints, demonstrating that our approach effectively synthesizes a traversable, 360° viewable scene.
-More rendered videos and navigable Gaussian Splats scenes can be found in the supplementary documents.
-A cottage in the woods where the walls are made of interwoven tree roots, and the floor is a carpet of moss.
-DreamScene360
-[55]
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Ours
-A cottage surrounded by fields of wildflowers that change color with the seasons, glowing faintly at twilight.
-DreamScene360
-[55]
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Ours
-(a) Panorama Reference
-(b) Novel View 1
-(c) Novel View 2
-(d) Novel View 3
-(e) Novel View 4
-Figure 4: Comparison of scene generation results.
-We compare our method with DreamScene360 [55] for text-based scene generation, as both approaches use panoramas as an intermediate representation for generating 360° viewable scenes.
-The first column presents the panorama reference: for DreamScene360, it is a single generated panorama from a text-to-image model, whereas in our results, it is the first frame of our generated 360° panorama video.
-While the previous work can produce a 360° viewable scene, its range of motion is highly limited.
-When novel viewpoints (Novel View 1-4) deviate significantly from the initial position, their GS scene exhibits severe blurriness and artifacts.
-Besides, our method achieves better alignment with the input prompt and produces the correct global structure of the scene.
-4 Results
-We demonstrate that our pipeline achieves state-of-the-art quality in both 360° video generation (Sec.
-4.1) and 3D scene generation (Sec.
-4.2), offering significantly improved visual consistency and view coverage.
-More visual examples, including videos and Gaussian Splats with an interactive viewer, can be found in the supplementary material.
-To quantitatively evaluate our method, we prompt ChatGPT [28] to generate a list of 320 descriptions of 3D scenes.
-We then use CLIP distance [34] to assess prompt adherence and Q-Align [46] to evaluate visual quality.
-Additionally, we perform ablation studies in Sec.
-4.3 to evaluate our design choices, including mixed data training and masked loss, to ensure the generation of high-quality videos suitable for reconstruction.
-4.1 Text-to-360°-Video Generation
-Our generation model can generate temporally consistent 360° panoramic videos, exceeding the quality of the state-of-the-art panoramic video generation models.
-We compare our 360° video generation results with 360DVD [42] in Fig.
-5.
+360DVD 0.7643 0.6704 0.538/0.661 89.7% 41.8%
+Ours 0.7387 0.8334 0.615/0.717 0% 94.6%
+Table 2.
+We quantitatively evaluate our method against several text-to-scene models, including (Chung et al., 2023), (Zhou et al., 2024), and (Yang et al., 2024).
+To assess prompt adherence and visual quality, we compute CLIP distance and Q-Align metrics on the generated 360° images or frames across 320 test prompts randomly generated by ChatGPT.
+Note that LucidDreamer is an image-to-scene generation model which requires both text and image as inputs, and we generate the initial image using ChatGPT; therefore CLIP and Q-Align score evaluating the panorama synthesis quality is not applicable for LucidDreamer.
+Our method outperforms all baselines across these evaluations.
+Furthermore, we conduct a 4AFC user study to assess three key aspects of scene quality: Realism/Completeness/Traversal Consistency and report the selection percentages.
+[…]
+LucidDreamer N/A N/A 21.0% 10.0% 10.0%
+DreamScene360 0.8116 0.8074 7.0% 6.0% 6.0%
+LayerPano3D 0.7977 0.8312 34.0% 25.0% 29.0%
+Ours 0.7387 0.8334 38.0% 59.0% 55.0%
+4.1. Text-to-360°-Video Generation
+A 360° video generator that accurately reproduces the distribution of real-world video captures is key to making it a sufficient proxy for 3D scene representation.
+Our generation model can generate spatial-temporal consistent 360° panoramic videos, exceeding the quality of the state-of-the-art panoramic video generation models.
+We compare our 360° video generation results with 360DVD (Wang et al., 2024) in Fig.
+3.
 We showcase four evenly sampled frames comparing 360DVD and our 360° panorama video generation model.
-Our model generates videos that closely match the input text prompt and produce richer scene details compared to 360DVD.
-The higher average CLIP distance and Q-Align metric for scenes generated from the 320 prompts in Table 1 further demonstrate the superior prompt alignment and visual quality of our method.
+Our model generates videos that closely match the input text prompt and produce more realistic scene details compared to 360DVD.
+In Table 1, we evaluate generation quality across multiple axes, where our generator consistently outperforms previous SOTA in common image/video generation quality metrics e.g., CLIP (Radford et al., 2021), Q-align (Wu et al., 2023), VBench (Huang et al., 2024).
 Additionally, videos generated by our model cover a broader scene span compared to 360DVD, which exhibits only slight rotations and translations of the viewpoint as indicated in the marked green box (Fig.
-5).
-More importantly, since the primary goal of 360° panoramic video synthesis is to generate a coherent 3D scene, the quality of 360DVD’s results prevent it from effectively accomplishing this task.
+3).
+More importantly, since the primary goal of 360° panoramic video synthesis is to generate a coherent 3D scene, the quality of 360DVD’s results prevents it from effectively accomplishing this task.
 Its outputs lack structural consistency across views, exhibiting continuous morphing and noticeable changes, which undermine the perception of a stable 3D environment.
@@
 3.3.
-We find that the failure rate of COLMAP reaches $89.7\%$, and even in cases where COLMAP succeeds, the match rate remains very low, with only $41.8\%$ images being successfully matched, making their results unsuitable for 3D reconstruction.
+As shown in Table 1, we find that the failure rate of COLMAP reaches $89.7\%$ (COLMAP-FR), and even in cases where COLMAP succeeds, the match rate (COLMAP-MR) remains very low, with only $41.8\%$ images being successfully matched, making their results unsuitable for 3D reconstruction.
 In contrast, our model preserves view-consistent local structures across frames.
 We do not encounter any failure cases when running COLMAP, and the average image match rate is as high as $94.6\%$ across our test prompts.
-This is the key reason we achieve a well-reconstructed 3D scene.
-Ours DreamScene360 [55] LucidDreamer [10]
-View Coverage 360° 360° Wide-angle front view
-Movable Range Large Very Limited Limited
-Table 2: Our method achieves state-of-the-art view coverage and traversable range compared to previous 3D generation methods.
-4.2 Text-to-Scene Generation
+This is the key reason our 360° videos can be a reasonable 3D representation for reconstruction.
+4.2. Text-to-Scene Generation
 As shown in Fig.
-3, by chaining our 360° video generator with the 3D reconstruction process, we enable the creation of a diverse set of 3D scenes, allowing users to navigate the environment from novel viewpoints and spatial locations.
+6, by chaining our 360° video generator with the 3D reconstruction process, we enable the creation of a diverse set of 3D scenes, allowing users to navigate the environment from novel viewpoints and spatial locations.
 More results can be found in the supplementary material.
 In Fig.
-4, we compare our results to the previous state-of-the-art text-to-scene method that can also create 360° virtual environment.
-However, as an approach that reconstructs a scene from only a single panorama, DreamScene360 cannot create a truly navigable scene.
-When synthesizing novel views (Novel View 1-4 in Fig.
-4) from locations different from the initial position, the rendered images exhibit significant artifacts and blurriness.
-In Table 1, we compute the Q-align metrics on the rendered novel views (Q-align NVS), where our results achieve higher visual quality and less blurriness compared to previous work, demonstrating that our generated scene enables a more realistic navigation experience.
-Similarly, we also report the CLIP distance and Q-Align metrics on the generated panorama images, demonstrating that our method generates a 3D environment that more accurately responds to user requests while achieving superior visual quality, even at the initial position.
-In addition to quantitative measurements, we highlight that our method is the first to achieve traversable and 360° viewable text-to-scene generation.
-As shown in Table 2, compared to previous work, our method achieves a complete 360° viewing area while also allowing users to walk inside the generated 3D scene for a large range.
-4.3 Ablation Study
+7, we compare our results to the previous state-of-the-art text-to-scene method that can also create 3D scenes.
+LucidDreamer (Chung et al., 2023), an outpainting-based scene generation method, synthesizes single-perspective views using an image diffusion model, but suffers from global geometric distortions and lacks support for full-view navigational experiences.
+Other methods that construct 3D scenes from panoramic images—such as DreamScene360 (Zhou et al., 2024) and LayerPano3D (Yang et al., 2024)—provide 360° view coverage, but do not enable truly navigable scenes, as significant camera movement often breaks scene coherence.
+For all other methods, when synthesizing novel views (Novel View 1–4 in Fig.
+7) from positions different from the initial viewpoint, the rendered images exhibit significant artifacts, including blurriness, missing regions, and structural distortions.
+Similarly, we report the CLIP distance and Q-Align metrics on the generated 360° images, demonstrating that our method generates a 3D environment that more accurately responds to user requests while achieving superior visual quality, even at the initial position.
+Figure 5.
+We verify that incorporating additional panorama image data helps the model generalize to diverse environments.
+Without the image data, the generated 360° videos not only exhibit degraded visual appearance but also display messy and incorrect content at the bottom of the video.
+This issue arises because these pixels lack supervision due to the masked loss.
+We argue that quality assessment is best conducted within the context of a real navigation experience.
+To this end, we perform a user study in which participants are presented with navigation videos generated by different methods.
+For each approach, we render a short video clip simulating camera movement through the scene, and ask users to select the most preferred result.
+Participants are instructed to evaluate the methods using a four-alternative forced-choice (4AFC) paradigm for the following dimensions: (1) Realism: Realistic Visual (2) Completeness: No missing regions (3) Traversal Consistency: Stable, distortion-free geometry during navigation.
+We collect 20 responses and report the selection percentages in Table 2.
+Our method received the majority of votes across all evaluation criteria, with particularly strong margins in Completeness and Traversal Consistency.
+4.3. Ablation Study
 Training a high-quality 360° video generator within our 3D generative pipeline is a key factor in the success of our method, relying on the use of mixed training data and the masked diffusion loss.
 We conduct ablation studies on these design choices to demonstrate their effectiveness.
-4.3.1 Mixed Training Data
+4.3.1. Mixed Training Data
 We conduct an experiment by excluding the additional image panorama data.
 As shown in Fig.
-6, incorporating image data significantly enhances the generalization capability of the video generator across a wide variety of scenes, and improves visual quality.
+5, incorporating image data significantly enhances the generalization capability of the video generator across a wide variety of scenes, and improves visual quality.
 Notably, due to the masked diffusion loss training, ground-truth pixels in the bottom regions of the panoramic videos are unavailable.
-The image data compensate for this limitation, enabling the model to intelligently auto-inpaint the bottom regions, resulting in a natural and realistic appearance.
-4.3.2 Masked Diffusion Loss
+The image data compensates for this limitation, enabling the model to intelligently auto-inpaint the bottom regions, resulting in a natural and realistic appearance.
+4.3.2. Masked Diffusion Loss
 We also train a model using the full diffusion loss applied to all pixels.
 The generated videos in this case reveal the photographer and the camera pod, which consistently appear in the 360° video training data (Fig.
-7).
+4).
 This obstructs a significant portion of the generated capture videos, making comprehensive scene reconstruction challenging.
 By applying the masked diffusion loss and incorporating the aforementioned image panorama data, we guide the model to generate clean 360° videos that can be directly used for perspective cropping and reconstruction.
-A cozy cabin library overlooking a vast snowy canyon.
-360DVD [42] Refer to caption Refer to caption Refer to caption Refer to caption
-Ours
-A coastal castle perched on a cliff under a dramatic sunset.
-360DVD [42] Refer to caption Refer to caption Refer to caption Refer to caption
-Ours
-Figure 5: Comparison of 360° video generation results.
-We present a comparison of 360° panoramic video generation results against 360DVD [42].
-Our generated videos exhibit significantly better visual quality and prompt alignment.
-We manually mark an anchor region with a green box in the 360° image, highlighting the substantially longer traversal range of our generated videos.
-A cozy cottage surrounded by wildflowers that glow in the moonlight, nestled deep in an enchanted forest.
-w/o Image Data Refer to caption Refer to caption
-w/ Image Data
-A serene desert where the sand shifts in swirling patterns, reflecting the colors of the sunset.
-w/o Image Data Refer to caption Refer to caption
-w/ Image Data
-Figure 6: We verify that incorporating additional panorama image data helps the model generalize to diverse environments.
-Without the image data, the generated 360° videos not only exhibit degraded visual appearance but also display messy and incorrect content at the bottom of the video.
-This issue arises because these pixels lack supervision due to the masked loss.
-A garden where every flower blooms in perfect circles, their petals reflecting the colors of the stars above.
-Baseline Refer to caption Refer to caption
-w/ Masked Loss
-A room where the ceiling is made of glass, revealing the swirling, ever-changing night sky beyond.
-Baseline Refer to caption Refer to caption
-w/ Masked Loss
-Figure 7: We conduct an ablation study to evaluate the impact of the masked diffusion loss.
-This loss function effectively guides the model to generate only the pixels of the scene, eliminating the moving photographer present in the original video training data.
-As a result, it enables the generation of a complete 360° scene capture for reconstruction.
-5 Conclusion
-In conclusion, we presented WorldPrompter, the first generative pipeline capable of producing a navigable 3D scenes from text prompts.
-By leveraging 360° panoramic video generation as an intermediate representation and a fast 3D reconstruction model, WorldPrompter enables the synthesis of immersive environments with comprehensive view coverage and high visual quality.
-This approach bridges the gap between text-to-3D generation and full scene synthesis.
-Although WorldPrompter achieves state-of-the-art performance, generating longer panoramic videos, spanning minutes rather than seconds, remains an open challenge.
-Consistent content creation over extended sequences is crucial for scaling up to larger, more detailed 3D scenes.
-Addressing this challenge would further expand applications in virtual worlds and beyond, paving the way for future advancements in scene-level generative models.
+4.4. Discussions and Limitations
+The quality of our text-to-scene generation pipeline is constrained by both the length and resolution of the panoramic video generative model.
+On one hand, generating minute-long panoramic videos while maintaining visual consistency across extended sequences is essential for scaling to larger, more detailed 3D scenes.
+On the other hand, panoramic videos allocate more pixels to capture global 3D content and ensure spatial consistency, but this comes at the cost of fine-grained details compared to perspective videos at the same resolution.
+We address this limitation with an upsampler, though a more fundamental solution would be a high-resolution video generator.
+5. Conclusion
+In conclusion, we demonstrate that 360° video capture is an effective and practical representation for scene-level 3D generation, as validated by WorldPrompter, a generative pipeline capable of producing navigable 3D scenes from text prompts.
+By leveraging 360° panoramic video generation as an intermediate representation, combined with a fast 3D reconstruction module, the pipeline enables the synthesis of immersive environments with comprehensive view coverage and high visual fidelity.
+This scalable approach effectively bridges the gap between text-to-3D generation and full-scene synthesis.
+We expect this novel idea highlights a promising research direction and has the potential to inspire future work in the field.
+Figure 6.
+Scene generation results.
+We present our scene generation results by showing (1) the first 360° image from our generated panoramic video as reference, and (2) novel video renderings (Novel View 1-4) from different spatial locations and viewpoints, demonstrating that our approach effectively synthesizes a traversable, 360° viewable scene.
+More rendered videos can be found in the supplementary documents.
+Figure 7.
+Comparison of scene generation results.
+We compare our method with LucidDreamer (Chung et al., 2023), DreamScene360 (Zhou et al., 2024) and LayerPano3D (Yang et al., 2024) for text-based scene generation.
+LucidDreamer generates scenes progressively from a perspective images, synthesizing views one at a time.
+DreamScene360 and LayerPano3D both use a single panorama as input to an image diffusion model, shown in the first column for reference.
+For our method, we present the first frame of the generated 360° panoramic video as the reference image.
+While prior methods can produce visually reasonable results at the initial viewpoint, their range of motion is limited.
+When synthesizing novel viewpoints (Novel View 1–4) that deviate significantly from the starting position, their reconstructed Gaussian Splatting (GS) scenes exhibit severe artifacts, including blurriness, missing regions, and distorted geometries.
+As a result, global structural consistency cannot be held.
+In contrast, our method produces scenes that remain well-aligned with the input prompt and maintain high visual quality even at viewpoints far from the initial position.